the interior angles of a triangle measures 2x,x+15 and 2x+15. what is the value of x?
Answer by amarjeeth123(574)

The interior angles of a triangle are 2x,x+15 and 2x+15.
We now that the sum of angles of a triangle is 180 degrees.
Plugging in the values we get,
2x+x+15+2x+15=180
5x+30=180
5x=180-30
5x=150
x=30
The value of x is 30.
Answer=30